# HiFo-Prompt: Prompting with Hindsight and Foresight for LLM-based Automatic Heuristic Design

### v1

@@ 3.3 Foresight: The Evolutionary Navigator for State-Aware Guidance
While the Hindsight module grounds the search process in historically validated principles, we propose the Foresight module, an Evolutionary Navigator that provides real-time, state-aware guidance for heuristic design. Its primary role is to dynamically regulate the balance between exploration and exploitation throughout the evolutionary process. It manages the evolutionary strategy by implementing a control policy $\pi$ that maps the macroscopic evolutionary state, $S_{t}$, to a high-level strategic orientation known as the Evolutionary Regime, defined as follows:
$\theta_{t}=\pi(S_{t})$ (6)
The control policy $\pi$ employs a rule-based logic to determine $\theta_{t}$, prioritizing responses to critical evolutionary dynamics such as performance stagnation or diminishing diversity. The macroscopic state $S_{t}$ encapsulates these dynamics through key metrics, formalized as:
@@
$\theta_{t}=\begin{cases}\theta_{\text{explore}}&\text{if }C_{\text{stag}}(t)\geq T_{\text{stag}}\text{ or }\Delta_{p}(t)<\Delta_{\text{crit}}\\ \theta_{\text{exploit}}&\text{if }C_{\text{prog}}(t)\geq T_{\text{prog}}\\ \theta_{\text{balance}}&\text{otherwise}\end{cases}$ (7)
where $C_{\text{stag}}(t)$ and $C_{\text{prog}}(t)$ are counters for consecutive generations of stagnation and progress, respectively, and $\Delta_{p}(t)$ measures population diversity. Specifically, stagnation is defined as a lack of improvement in the best-so-far fitness, while progress requires a significant fitness gain. Population diversity, $\Delta_{p}(t)$, is quantified as the standard deviation of the fitness values within the current population.
The hyperparameters $T_{\text{stag}}$, $T_{\text{prog}}$, and $\Delta_{\text{crit}}$ are thresholds that ensure the Navigator is highly responsive to the evolving search landscape. The thresholds for stagnation ($T_{\text{stag}}$) and progress ($T_{\text{prog}}$) are intentionally set to small values, a design choice motivated by the rapid convergence of our framework. Unlike traditional Evolutionary Algorithms that may evolve over many generations, HiFo-Prompt often achieves significant progress in very few iterations, necessitating a Navigator that can react swiftly to these short-term dynamics. Similarly, the critical diversity threshold, $\Delta_{\text{crit}}$, is carefully calibrated to balance the need to maintain population diversity to converge on high-quality solutions.
Once the strategic regime is determined, the Evolutionary Navigator translates this strategy into a concrete instruction for the LLM, known as a Design Directive. The nature of this directive is dictated by the active regime ($\theta_{t}$). For example, the $\theta_{\text{explore}}$ regime yields directives that prompt radical innovation, such as invent a new crossover operator from scratch. Conversely, the $\theta_{\text{exploit}}$ regime issues directives focused on incremental refinement, like slightly modify the selection pressure of the best heuristic. The $\theta_{\text{balance}}$ regime generates directives that synthesize these approaches, such as combine the strengths of the top two performing heuristics. This mechanism ensures that the generative focus of the LLM is dynamically steered to align with the strategic intent of the current evolutionary phase.
@@ 4 Experiments
Method TSP10 TSP20 TSP50 TSP100 TSP200
Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj.
LKH3 0.000% 2.846 0.000% 3.863 0.000% 5.709 0.000% 7.793 0.000% 10.727
EoH 6.480% 3.030 9.236% 4.220 11.553% 6.369 13.638% 8.856 14.209% 12.251
ReEvo 5.215% 2.994 6.156% 4.101 9.714% 6.264 11.379% 8.680 12.830% 12.103
HSEvo 5.075% 2.990 7.320% 4.146 11.185% 6.348 11.944% 8.724 13.463% 12.171
MCTS-AHD 4.829% 2.983 8.045% 4.174 10.642% 6.317 12.521% 8.769 13.510% 12.176
Ours 1.560% 2.890 2.790% 3.971 5.808% 6.041 8.158% 8.429 9.315% 11.726
Table 1: Results on TSP with step-by-step construction. Gap(%) denotes the performance gap compared to advanced heuristic algorithms. Obj. represents the value of the objective function. The best results are highlighted in bold.
@@
POMO 0.189% 0.496% 2.935%
LEHD 0.099% 0.131% 0.342%
EoH 0.000% 0.026% 0.453%
HSEvo 0.000% 0.030% 0.821%
Ours 0.000% 0.014% 0.397%
Table 2: Results on TSP with GLS. Comparison relative to the results of advanced heuristic on TSP50, TSP100 and TSP200. The best-performing overall and LLM-based AHD results are shown in bold.
In this section, we present the results of heuristics designed by our proposed HiFo-Prompt on different complex tasks, including Traveling Salesman Problem (TSP)(Matai2010, ), Online Bin Packing Problem (Online BPP) (Seiden2002Online, ), Flow Shop Scheduling Problem (FSSP) (Emmons2012FlowShop, ), and Bayesian Optimization (BO) (Shahriari2016Taking, ). Task definitions and details are given in Appendix B. Results on TSP, online BPP, and FSSP are presented in this section, while the results on BO are provided in Appendix C.4, where HiFo-Prompt demonstrates competitive and reliable performance.
Settings.
We use the qwen2.5-max model via the DashScope API with a temperature of 1.0 and evolve a population of 4 individuals. The Insight Pool in Hindsight module has a capacity of $C_{\text{pool}}=30$, with a Jaccard similarity threshold of $0.7$ for deduplication. Insight retrieval uses a utility function with selection count $s=3$, usage penalty weight $w_{u}=0.1$, and recency bonus $\tau_{r}=0.2$. Credit assignment is performed using EMA with a learning rate of $\alpha=0.3$. For pool maintenance, each insight’s eviction score decays at $R_{\text{decay}}=0.01$ per generation, and new insights receive a grace period of $T_{\text{usage}}=3$ applications. In Foresight module, Evolutionary Navigator’s thresholds are set to $T_{\text{stag}}=3$ for stagnation, $T_{\text{prog}}=2$ for progress, and $\Delta_{\text{crit}}=0.3$ for critical diversity. We run 8 generations for CO tasks and 4 for BO, where rapid convergence demonstrates our method’s efficiency. To ensure fairness, all LLM-based AHD baselines use the same LLM endpoint and query budget, with other settings kept as reported in their original papers.
Baselines.
To demonstrate the effectiveness of our proposed method in designing heuristics, we introduce several approaches for solving these complex optimization tasks. (1) handcrafted heuristics, e.g., LKH3(LinKernighan1973, ) for TSP, First Fit and Best Fit (romera2024, ) for Online BPP, NEH (nawaz1983heuristic, ) and NEHFF (FernandezViagas2014, ) for FSSP. (2) Neural Combinatorial Optimization (NCO) methods, e.g., POMO (kwon2020pomo, ) and LEHD (luo2023neural, ) for TSP, PFSPNet_NEH (pan2022deep, ) for FSSP. (3) LLM-based AHD methods, e.g., Funsearch (romera2024, ), EoH (liu_icml2024, ), ReEvo (ye2024, ), HSEvo (Dat2025HSEvo, ) and MCTS-AHD (zheng2025, ). Most of our test datasets follow EoH. Notably, Funsearch, ReEvo, and HSEvo require a seed function to initialize their populations, while EOH, MCTS-AHD, and our method can run without it.
@@
First Fit* 4.40% 4.18% 4.27%
Best Fit* 4.08% 3.83% 3.91%
Funsearch* 0.80% 1.07% 1.47%
EOH 1.07% 1.07% 1.06%
ReEvo 0.78% 4.47% 3.24%
HSEvo 0.79% 5.41% 4.04%
MCTS-AHD 0.99% 0.95% 0.95%
Ours 0.66% 0.63% 0.66%
Table 3: Results on Online BPP. The results denote the ratio of excess bins compared to the lower bound on Weibull instances. Results with * are from EoH (liu_icml2024, ). The best results are highlighted in bold.
Traveling Salesman Problem.
We design heuristics for solving the TSP using both Step-by-Step Construction (asani2023convex, ) and Guided Local Search (GLS) (alsheddy2016guided, ). Table 1 compares the results of our method with other baselines. We evaluated the performance on 100 instances at each of five sizes. To demonstrate performance on out-of-distribution instances, we also conducted experiments on the TSPLib dataset (reinelt1991tsplib, ), and the results can be found in the Appendix C.1. We further evaluated our method on 100 instances of TSP50, TSP100, and TSP200 in GLS. The results are shown in Table 2. The heuristic designed by our method achieves a significant performance improvement compared to the LLM-based AHD approach. It can even outperform or achieve comparable performance to some outstanding NCO methods (vinyals2015pointer, ) on small-scale instances.
nn20,mm10 nn50,mm10 nn100,mm10
NEH* 4.05% 3.47% 2.07%
NEHFF* 4.15% 3.62% 1.88%
PFSPNet_NEH* 4.04% 3.48% 1.72%
EOH 0.244% 0.287% 0.174%
Ours 0.145% 0.120% 0.094%
nn20,mm20 nn50,mm20 nn100,mm20
NEH* 3.06% 5.48% 3.58%
NEHFF* 2.72% 5.10% 3.73%
PFSPNet_NEH* 2.96% 5.05% 3.56%
EOH 0.179% 0.753% 0.927%
Ours 0.103% 0.600% 0.472%
Table 4: Results on FSSP. The results show the comparison of makespan relative to the baseline on Taillard instances. $n$ denotes the number of jobs and $m$ denotes the number of machines. Results with * are from EoH (liu_icml2024, ). The best results are highlighted in bold.
@@
HiFo-Prompt 1.99% 0.66% 2.01% 0.63% 1.95% 0.66%
Online Bin Packing Problem.
The key function of Online BPP is a scoring function that outputs a score for each bin, based on the current item’s size and the remaining capacities of the bins. We evaluate our method on Weibull BPP instances (romera2024, ) following the EoH setting. The results with three scales are shown in Table 3. Our method not only significantly outperforms handcrafted heuristics but also surpasses LLM-based AHD approaches. More results are provided in Appendix C.2.
Flow Shop Scheduling Problem.
FSSP involves scheduling $n$ jobs on $m$ machines to minimize the makespan, where each job comprises $m$ operations processed in a fixed order. We evaluate the heuristic designed by our method on Taillard instances (taillard1993benchmarks, ). The dataset includes instances with 20 to 100 jobs ($n$) and 10 to 20 machines ($m$). Table 4 presents some of the results, with additional results provided in Appendix C.3. Our method performs well on all datasets, consistently delivering strong and reliable results across different scenarios.
Figure 3: Convergence Analysis for TSP and Online BPP.
@@
Comparative Results.
We compared the progression of objective function values during the evolutionary process of our method and EoH on both the TSP and Online BPP. Figure 3 presents the convergence analysis curves. The figure demonstrates that our method converges more rapidly while requiring fewer individuals in the population, indicating higher efficiency. In addition, to further highlight the efficiency of our method in terms of training time and computational cost, we measured both the training duration and the number of LLM requests for the four methods. The results are presented in Figure 4.
Figure 4: Efficiency Comparison on LLM Requests and Training Time for TSP and Online BPP.
@@ A.1 Problem Formulation of Automatic Heuristic Design
where $g(h)$ is a fitness function that maps a heuristic to a real number, estimating its quality based on its expected performance on a representative set of problem instances $D$. For a task with a minimization objective $f$ (e.g., minimizing cost or error), this performance metric is defined as:
@@ A.2 LLM-driven Evolutionary Computation
This framework, introduced within the "heuristic evolution" paradigm, leverages LLM-driven generation to explore the heuristic space and a selection mechanism to exploit high-performing solutions.

### v2

@@ 3.3 Foresight: The Evolutionary Navigator for State-Aware Guidance
While the Hindsight module grounds the search in historically validated principles, the Foresight module acts as an Evolutionary Navigator, providing real-time, state-aware guidance. Its primary role is to modulate the exploration-exploitation trade-off via a control policy, $\pi:S_{t}\rightarrow\theta_{t}$. This policy maps the current evolutionary state $S_{t}$ to a high-level search orientation, denoted as the Evolutionary Regime $\theta_{t}$. Unlike traditional adaptive EAs that solely tune numerical parameters, Foresight establishes a semantic feedback loop that directly steers the conceptual strategy of the LLMs. We define $\pi$ as a rule-based system that selects the appropriate strategic modes based on evolutionary state :
@@
$\theta_{t}=\begin{cases}\theta_{\text{explore}}&\text{if }C_{\text{stag}}(t)\geq\tau_{\text{stag}}\text{ or }\Delta_{p}(t)<\delta_{p}\\ \theta_{\text{exploit}}&\text{if }C_{\text{prog}}(t)\geq\tau_{\text{prog}}\\ \theta_{\text{balance}}&\text{otherwise}\end{cases}$ (6)
To implement this policy, the Navigator continuously monitors two primary aspects of the evolutionary search: its performance trajectory and its population diversity. The performance trajectory is tracked via two mutually exclusive counters, $C_{\text{prog}}$ and $C_{\text{stag}}$, based on the improvement in the best raw fitness $\Delta g$. If $\Delta g>10^{-4}$, it is a generation of progress (increment $C_{\text{prog}}$, reset $C_{\text{stag}}$); otherwise, it is stagnation (increment $C_{\text{stag}}$, reset $C_{\text{prog}}$). Simultaneously, the Navigator assesses population health through a novel measure of phenotypic diversity, $\Delta_{p}(t)$. This metric is crucial because relying on fitness alone can be misleading; a population of high-fitness but structurally similar individuals often indicates entrapment in a local optimum. To counteract this, our metric quantifies the semantic variety within the population by measuring the dissimilarity of the generated algorithms’ textual forms, rather than their fitness values. It is computed as the normalized fraction of unique pairs of algorithms in the population $P$ whose textual descriptions are non-identical:
$\Delta_{p}(t)=\frac{1}{|P|(|P|-1)/2}\sum_{i=1}^{|P|}\sum_{j=i+1}^{|P|}\mathbb{I}(\text{alg}_{i}\neq\text{alg}_{j})$ (7)
where $\mathbb{I}(\cdot)$ is the indicator function and $\text{alg}_{i}$ denotes the textual description of the $i$-th algorithm. Unlike embedding-based metrics, which risk obscuring critical logical alterations (e.g., dynamic vs. static) through semantic smoothing (Agarwal et al., 2025), this metric relies on exact string matching of the generated algorithmic descriptions. By enforcing a standardized prompt template to eliminate trivial syntactic noise, any remaining lexical difference indicates a deliberate structural modification by the LLM. Consequently, this approach provides a robust and computationally efficient proxy for phenotypic diversity, ensuring that even subtle yet functional changes are preserved.
This hybrid state representation, which marries quantitative performance trends with a qualitative measure of semantic diversity, provides the Navigator with a much more holistic understanding of the search landscape than fitness-based metrics alone can. These state indicators are evaluated against empirically determined thresholds, with their specific values detailed in our experimental setup. The progress and stagnation counters are designed to track immediate performance trends. This focus on recent history is crucial for capturing the rapid dynamics inherent in LLM-based evolution, enabling swift strategic adjustments in response to even short periods of stagnation or consistent improvement. Similarly, the diversity threshold is calibrated to detect a significant collapse in semantic variety. It acts as an early warning mechanism against premature convergence, triggering stronger exploratory pressure when a substantial portion of the population becomes homogeneous. Once the regime $\theta_{t}$ is determined, it is translated into a natural language Design Directive(See Section F.2) that is injected into the LLM’s prompt. For instance, $\theta_{\text{explore}}$ might yield a directive to try a significantly different approach from conventional solutions, whereas $\theta_{\text{exploit}}$ would instruct the model to focus on refining and optimizing the most effective patterns. This mechanism ensures the LLM’s generative focus is explicitly aligned with the high-level strategy dictated by the current evolutionary state.
@@ 4 Experiments
In this section, we present the results of heuristics designed by our proposed HiFo-Prompt on different complex tasks, including Traveling Salesman Problem (TSP) (Matai et al., 2010), Online Bin Packing Problem (Online BPP) (Seiden, 2002), Flow Shop Scheduling Problem (FSSP) (Emmons and Vairaktarakis, 2012), and Bayesian Optimization (BO) (Shahriari et al., 2016). Task definitions and details are given in Appendix B. Results on TSP, online BPP, and FSSP are presented in this section, while the results on BO are provided in Appendix C.4, where HiFo-Prompt demonstrates competitive and reliable performance.
Experimental Settings.
Experiments were conducted on a workstation equipped with an Intel Core i7-12700 CPU, employing Qwen2.5-Max as the core LLM. The evolutionary process maintains a population size of 8, running for 8 generations on Combinatorial Optimization (CO) tasks and 4 generations on Bayesian Optimization (BO) tasks. Regarding specific module hyperparameters, the Hindsight Module maintains an Insight Pool of capacity $C_{\text{pool}}=30$ (Jaccard threshold $0.7$), with retrieval parameters set to selection count $s=3$, usage penalty $w_{u}=0.1$, recency bonus magnitude $\tau_{r}=0.2$, EMA rate $\alpha=0.3$, decay $R_{\text{decay}}=0.01$, and probation $T_{\text{usage}}=3$. The Foresight Module utilizes Navigator thresholds of $\tau_{\text{stag}}=3$ (stagnation), $\tau_{\text{prog}}=2$ (progress), and $\delta_{p}=0.3$ (diversity). To ensure fair comparison, all LLM-based AHD baselines utilize the same underlying LLM, while their hyperparameters follow the settings reported in their original publications.
Baselines.
To demonstrate the effectiveness of our proposed method in designing heuristics, we introduce several approaches for solving these complex optimization tasks. (1) handcrafted heuristics, e.g., LKH3(Lin and Kernighan, 1973) for TSP, First Fit and Best Fit (Romera-Paredes et al., 2024) for Online BPP, NEH (Nawaz et al., 1983) and NEHFF (Fernandez-Viagas and Framinan, 2014) for FSSP. (2) Neural Combinatorial Optimization (NCO) methods, e.g., POMO (Kwon et al., 2020) and LEHD (Luo et al., 2023) for TSP, PFSPNet_NEH (Pan et al., 2022) for FSSP. (3) LLM-based AHD methods, e.g., Funsearch (Romera-Paredes et al., 2024), EoH (Liu et al., 2024b), ReEvo (Ye et al., 2024), HSEvo (Dat et al., 2025) and MCTS-AHD (Zheng et al., 2025). Most of our test datasets follow EoH. Notably, Funsearch, ReEvo, and HSEvo require a seed function to initialize their populations, while EoH, MCTS-AHD, and our method can run without it.
Traveling Salesman Problem.
Step-by-Step Construction (Asani et al., 2023) and Guided Local Search (GLS) (Alsheddy et al., 2016) are two different strategies for solving TSP. The detailed procedure of the two strategies can be found in the Appendix B.1. We design key heuristics for these two strategies. Table 1 compares the results of our method with other baselines in step-by-step construction. We evaluated the performance on 100 instances at each of three sizes. To demonstrate performance on out-of-distribution instances, we also conducted experiments on the TSPLib dataset (Reinelt, 1991), and the results can be found in the Appendix C.1. We further evaluated our method on 100 instances of TSP100, TSP200, and TSP500 in GLS. The results are shown in Table 2. The heuristic designed by our method achieves a significant performance improvement compared to LLM-based AHD methods.
Table 1: Results on TSP with step-by-step construction. Gap(%) denotes the performance gap compared to advanced heuristic algorithms. Time(s) represents the running time of the designed heuristics. This result of LLM-based AHD method is the average of three runs. The best-performing LLM-based AHD results are shown in bold.
@@
Gap Time(s) Gap Time(s) Gap Time(s)
LKH3 0.000% 323.3 0.000% 1450 0.000% 6312
POMO 0.163% - 1.636% - 13.961% -
LEHD 0.117% - 0.452% - 0.367% -
EoH 12.820% 1.4 15.361% 9 16.658% 78
ReEvo 10.239% 21.5 12.577% 224 14.890% 3013
HSEvo 10.467% 89.5 12.008% 1286 13.578% 24835
MCTS-AHD 10.642% 91.5 12.521% 1084 13.510% 14521
Ours 6.625% 244.7 8.582% 1843 8.877% 16099
Table 2: Results on TSP with GLS. Comparison relative to the results of advanced heuristic on TSP100, TSP200 and TSP500. The result of the LLM-based AHD method is the average of three runs. The best results are shown in bold.
Method TSP100 TSP200 TSP500
Gap Time(s) Gap Time(s) Gap Time(s)
LKH3 0.000% - 0.000% - 0.000% -
EoH 0.026% 210.3 0.453% 368.1 2.037% 1100.7
ReEvo 0.049% 357.1 0.424% 775.8 2.090% 1103.4
HSEvo 0.087% 543.4 0.886% 792.9 2.507% 1105.1
Ours 0.015% 217.0 0.382% 392.4 1.520% 1100.8
Table 3: Results on Online BPP. Gap(%) denotes the ratio of excess bins compared to the lower bound on Weibull instances. Obj. represents the value of the objective function. Results with * are from EoH (Liu et al., 2024b). This result of LLM-based AHD method is the average of three runs. The best results are highlighted in bold.
@@
Gap Obj Gap Obj Gap Obj
lower bound 0.00% 2006.2 0.00% 1740.2 0.00% 1687.0
First Fit* 4.40% - 4.18% - 4.27% -
Best Fit* 4.08% - 3.83% - 3.91% -
Funsearch* 0.80% 2022.2 1.07% 1758.8 1.47% 1711.8
EoH 1.02% 2026.6 1.00% 1757.6 1.00% 1703.9
ReEvo 0.78% 2021.8 4.47% 1818.0 3.24% 1741.6
HSEvo 1.91% 2044.6 5.47% 1835.4 4.39% 1761.1
MCTS-AHD 0.99% 2026.0 0.95% 1756.8 0.95% 1703.0
Ours 0.69% 2020.1 0.66% 1751.7 0.66% 1698.1
Online Bin Packing Problem.
The key function of Online BPP is a scoring function that outputs a score for each bin, based on the current item’s size and the remaining capacities of the bins. We evaluate our method on Weibull BPP instances (Romera-Paredes et al., 2024) following the EoH setting. The results with three scales are shown in Table 3. Our method not only significantly outperforms handcrafted heuristics but also surpasses LLM-based AHD approaches. More results are provided in Appendix C.2.
Flow Shop Scheduling Problem.
FSSP involves scheduling $n$ jobs on $m$ machines to minimize the makespan, where each job comprises $m$ operations processed in a fixed order. We evaluate the heuristic designed by our method on Taillard instances (Taillard, 1993). The dataset includes instances with 20 to 100 jobs ($n$) and 10 to 20 machines ($m$). Table 4 presents some of the results, with additional results provided in Appendix C.3. Our method performs well on all datasets, consistently delivering strong and reliable results across different scenarios.
Table 4: Results on FSSP. The results show the comparison of makespan relative to the baseline on Taillard instances. $n$ denotes the number of jobs and $m$ denotes the number of machines. Results with * are from EoH (Liu et al., 2024b). This result of LLM-based AHD method is the average of three runs. The best results are highlighted in bold.
Method n20,m10 n50,m10 n100,m10 n20,m20 n50,m20 n100,m20
NEH* 4.05% 3.47% 2.07% 3.06% 5.48% 3.58%
NEHFF* 4.15% 3.62% 1.88% 2.72% 5.10% 3.73%
PFSPNet_NEH* 4.04% 3.48% 1.72% 2.96% 5.05% 3.56%
EoH 0.31% 0.29% 0.23% 0.20% 0.84% 0.94%
Ours 0.17% 0.17% 0.13% 0.10% 0.58% 0.51%
@@
HiFo-Prompt 2.19% 0.69% 2.08% 0.66% 2.07% 0.66%
@@
Discussion.
We provide more experiments and discussion in the appendix. In Appendix C.5, we present comparative studies between our method and EoH in terms of convergence speed. We also evaluate our method across different LLMs to further demonstrate its effectiveness and generality. In Appendix C.6, we conduct a thorough sensitivity analysis of the key hyperparameters involved in our method, providing detailed explanations of their impact on performance. In Appendix C.7, we conduct additional ablation studies and statistical analyses to further investigate the necessity, underlying mechanisms, and stability of the key design components of HiFo-Prompt.
@@ 5 Conclusion
Acknowledgments
This work was partially supported by National Natural Science Foundation of China (grants 12571590, 12426305, 62401473), National Key Research and Development Program of China (grant 2022YFA1004201), Tianyuan Fund for Mathematics of the National Natural Science Foundation of China (grant 12426105), Fundamental Research Funds for the Central Universities (grant xzd012024049), National Key Laboratory Fund Project for Space Microwave Communication (grant HTKJ2024KL504010), Shenzhen Science and Technology Program (grant JCYJ20240813150735045) and Key Research and Development Project in Shaanxi Province (grant 2025CY-YBXM-055).
@@ A.1 Problem Formulation of Automatic Heuristic Design
where $g(h)$ is a fitness function (maximized) that maps a heuristic to a real number, estimating its quality based on its expected performance on a representative set of problem instances $D$. For a task with a minimization objective $f$ (minimized, e.g., minimizing cost or error), this performance metric is defined as:
@@ A.2 LLM-driven Evolutionary Computation
This framework, introduced within the heuristic evolution paradigm, leverages LLM-driven generation to explore the heuristic space and a selection mechanism to exploit high-performing solutions.